# Elliptic boundary-value problems in some distribution spaces of generalized smoothness

### v1

@@ Abstract.
We build a solvability theory of elliptic boundary-value problems in normed Sobolev spaces of generalized smoothness for any integral-exponent $p>1$. The smoothness is given by a number parameter and a supplementary function parameter that varies slowly at infinity. These spaces are obtained by a combination of the methods of the complex interpolation with number parameter between Banach spaces and the quadratic interpolation with function parameter between Hilbert spaces applied to classical Sobolev spaces. We show that the spaces under study admit localization near a smooth boundary and describe their trace spaces in terms of Besov spaces with the same supplementary function parameter. We prove that a general differential elliptic problem induces Fredholm bounded operators on appropriate pairs of the spaces under study. We also find exact sufficient conditions for solutions of the problem to have a prescribed generalized or classical smoothness on a given set and establish corresponding a priory estimates of the solution. These results are specified for parameter-elliptic problems.
@@
Sobolev spaces; Besov spaces; generalized smoothness; interpolation functor; elliptic problem; parameter-elliptic problem; Fredholm operator; a priory estimate
@@ 1. Introduction
In the last two decades, the mathematical literature on function spaces demonstrates a growing interest in distribution spaces of generalized smoothness and their applications (see, e.g., [1, 7, 10, 11, 14, 21, 48, 50, 22, 35, 49]). These spaces characterize the smoothness properties of distributions by means of a function parameter depending on the frequency variables (or by a number sequence representing this function). This allows describing such properties far more finely than the scales of classical distribution spaces, which proved to be very useful in the approximation theory [59], theory of stochastic process [25], and theory of partial differential equations (PDEs) [44, 52]. As to boundary-value problems for PDEs, the applications of distributions spaces of generalized smoothness has been restricted only to Hilbert spaces [5, 6, 36, 37, 43, 44], which can be investigated with the help of Spectral Theorem applied to a self-adjoint operator (see, e.g., [48]). However, theorems on solvability of various classes of boundary-value problems are well known for non-Hilbert distribution spaces of classical smoothness given by a number (see, e.g., [2, 34, 38, 61, 62]). Concerning elliptic boundary-value problems, the expediency and prospects of their study in normed spaces of generalized smoothness are discussed by Triebel in [63, pp. 57–60].
@@
The purpose of this paper is to extend the classical theory of solvability of elliptic boundary-value problems in $L_{p}$-Sobolev spaces $H^{s}_{p}$ with $1<p<\infty$ (see, e.g., [2, Chapter V] and [62, Chapter 5]) to some Sobolev spaces of generalized smoothness. The latter spaces form the scale $\{H^{s,\varphi}_{p}\}$ calibrated with the help of a supplementary function parameter $\varphi$ much more finely than the Sobolev scale. The positive function $\varphi(t)$ of $t\geq 1$ varies slowly at infinity in the sense of Karamata, which implies that the space $H^{s,\varphi}_{p}$ is located between any spaces $H^{s-\varepsilon}_{p}$ and $H^{s+\varepsilon}_{p}$ with $\varepsilon>0$ on this scale. We study interpolation properties of the scale and show that every space $H^{s,\varphi}_{p}$ is obtained by the complex interpolation between certain spaces $H^{s_{0},\varphi_{0}}_{2}$ and $H^{s_{1}}_{p_{1}}$. Since the scale $\{H^{s,\varphi}_{2}\}$ of Hilbert spaces is obtained by the quadratic interpolation with function parameter between inner product Sobolev spaces [44, Theorem 1.14], we use these two interpolation methods to study the spaces $H^{s,\varphi}_{p}$ and elliptic problems in them. It is important for applications and follows from these interpolation properties that that the space $H^{s,\varphi}_{p}(\Omega)$ over a bounded Euclidean domain $\Omega$ with a boundary $\Gamma\in C^{\infty}$ admits localization and that the trace operator acts continuously form whole $H^{s,\varphi}_{p}(\Omega)$ onto the Besov space $B^{s-1/p,\varphi}_{p}(\Gamma)$ of the same supplementary generalized smoothness $\varphi$ whenever $s>1/p$. The distribution space $B^{\sigma,\varphi}_{p}(\Gamma)$ with arbitrary $\sigma\in\mathbb{R}$ also admits localization and has quite similar interpolation properties. These results are discussed in Section 2 of the paper.
@@
Section 4 gives an application of spaces (1.1) to a general parameter-elliptic boundary-value problem. We show that this problem sets an isomorphism between these spaces for sufficiently large absolute values of the parameter from the ellipticity angle of the problem. Moreover, the solutions of the problem admits a corresponding two-sided a priori estimate in parameter-dependent norms with constants that do not depend on the parameter. This result is based on interpolation properties of such norms. The results can be applied to the study of parabolic initial-boundary-value problems in some anisotropic distribution spaces of generalized smoothness (cf. [4, Chapter II]).
@@ 2. Spaces of generalized smoothness
belongs to $L_{p}(\mathbb{R}^{n})$. Here, as usual, $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is the linear topological space of all tempered distributions on $\mathbb{R}^{n}$; thus, $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is dual to the linear topological Schwartz space $\mathcal{S}(\mathbb{R}^{n})$ of all infinitely smooth rapidly decreasing functions on $\mathbb{R}^{n}$. Moreover, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform ant its inverse; $\langle\xi\rangle:=(1+|\xi|^{2})^{1/2}$ is the smoothed absolute value of the vector $\xi\in\mathbb{R}^{n}$, and $L_{p}(\mathbb{R}^{n})$ is the normed space of all Lebesgue measurable functions $h:\mathbb{R}^{n}\to\mathbb{C}$ such that $|h|^{p}$ is integrable over $\mathbb{R}^{n}$. The space $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ is endowed with the norm
@@
The space $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ belongs to the classes of Sobolev spaces of generalized smoothness introduced and investigated by Volevich and Paneach [64, § 13], Schechter [57, Section 4], Goldman [17], Triebel [60, Section 5], Merucci [39, Section 4], and others. This space is complete and separable and is continuously embedded in $\mathcal{S}^{\prime}(\mathbb{R}^{n})$, with $\mathcal{S}(\mathbb{R}^{n})$ being continuously and densely embedded in $H^{s,\varphi}_{p}(\mathbb{R}^{n})$. This follows from the analogous properties of $L_{p}(\mathbb{R}^{n})$ in view of the evident fact that the mapping $w\mapsto w_{s,\varphi}$ set an isometric isomorphism between $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ and $L_{p}(\mathbb{R}^{n})$. Hence, the set $C^{\infty}_{0}(\mathbb{R}^{n})$ of infinitely smooth compactly supported functions on $\mathbb{R}^{n}$ is dense in $H^{s,\varphi}_{p}(\mathbb{R}^{n})$.
If $\varphi(t)=1$ whenever $t\geq 1$, then $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ is the Sobolev space $H^{s}_{p}(\mathbb{R}^{n})$ with smoothness index $s$ and integral-exponent $p$. Generally, we have the dense continuous embeddings
@@
Indeed, it follows from property (i) that $t^{-\varepsilon/2}\varphi(t)\to 0$ and $t^{\varepsilon/2}\varphi(t)\to\infty$ as $t\to\infty$ [58, Section 1.5, property $1^{\circ}$]; hence, the functions $t^{s}\varphi(t)/t^{s+\varepsilon}$ and $t^{s-\varepsilon}/t^{s}\varphi(t)$ of $t\geq 1$ are integrable over $[1,\infty)$ with respect to $dt/t$, which implies (2.2) by [39, Subsection 4.1, Proposition 6].
@@
is naturally called the refined Sobolev scale (over $\mathbb{R}^{n}$) with integral-exponent $p$. In the $p=2$ case of Hilbert spaces, such a scale was introduced by Mikhailets and Murach [40] and applied to elliptic operators and elliptic boundary-value problems [41, 42, 43, 44, 51]. Various distribution spaces of generalized smoothness given by a slowly varying function were investigated by Haroske and Moura [20], specifically Besov spaces over $\mathbb{R}^{n-1}$ used below.
@@
$\{w\in H^{s,\varphi}_{p}(\mathbb{R}^{n}):\mathrm{supp}\,w\subset\overline{\Omega}\}.$
@@
If $p\neq 2$, then the refined Sobolev scale with integral-exponent $p$ is obtained by the complex interpolation between Hilbert spaces $H^{s_{0},\varphi_{0}}_{2}(\Omega)$ and Sobolev spaces $H^{s_{1}}_{p_{1}}(\Omega)$. Moreover, the collection of these scales is stable with respect to this interpolation. Let us formulate these properties in the form of two theorems. As usual, $[E_{0},E_{1}]_{\theta}$ denotes the result of the complex interpolation with the number parameter $\theta\in(0,1)$ between Banach spaces $E_{0}$ and $E_{1}$ continuously embedded in a Hausdorff topological linear space (see., e.g., [62, Section 1.9]). Such a pair of the spaces is called an interpolation pair.
@@
$[H^{s_{0},\varphi_{0}}_{p_{0}}(\Omega),H^{s_{1},\varphi_{1}}_{p_{1}}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$ (2.9)
@@
Let $s\in\mathbb{R}$, $\varphi\in\Upsilon$, and $p\in(1,\infty)$ with $p\neq 2$. We arbitrarily chose a real number $p_{1}$ such that $21<p_{1}<p<$ 2 or $2<p<p_{1}<\infty$ and define a number $\theta\in(0,1)$ by the formula
@@
Note that the interpolation formulas (2.7), (2.9), and (2.11) are especially useful when we study bounded linear operators on the corresponding spaces because the boundedness of such operators is preserved under the interpolation between these spaces. Namely, if $[E_{0},E_{1}]$ and $[G_{0},G_{1}]$ are interpolation pairs of Banach spaces and if a linear mapping $T$ is a bounded operator from $E_{j}$ to $G_{j}$ for each $j\in\{0,1\}$, then $T$ is a bounded operator from $[E_{0},E_{1}]_{\theta}$ to $[G_{0},G_{1}]_{\theta}$ whenever $10<\theta<$ 1. The quadratic interpolation between Hilbert spaces has an analogous property.
@@
+$\mathbb{R}^{n}_{+}:=\{(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}>0\},$
@@
+$\partial\mathbb{R}^{n}_{+}=\mathbb{R}^{n}_{0}:=\{(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}=0\in\mathbb{R}\}.$
@@
Choose functions $\chi_{1},\ldots,\chi_{\varkappa+r}\in C^{\infty}(\mathbb{R}^{n})$ such that $\mathrm{supp}\,\chi_{j}\subset U_{j}$ whenever $1\leq j\leq\varkappa+r$ and that $\chi_{1}+\cdots+\chi_{\varkappa+r}=1$ on $\overline{\Omega}$. Let $\mathcal{O}$ and $\mathcal{O_{+}}$ denote the operators of the extension of a function/distribution by zero onto $\mathbb{R}^{n}$ and +$\mathbb{R}^{n}_{+}$, resp.
@@
where $u\in\mathcal{S}\,^{\prime}(\Omega)$, sets a bounded linear operator
@@
where $v_{1},\ldots,v_{\varkappa}\in\mathcal{S}^{\prime}(\mathbb{R}^{n}_{+})$ and $v_{\varkappa+1},\ldots,v_{\varkappa+r}\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Here, $\mathcal{O}_{\Omega}$ denotes the operator of the extension of a distribution by zero onto $\Omega$. Then $KTu=u$ for every $u\in\mathcal{S}^{\prime}(\Omega)$. This mapping sets a bounded linear operator
@@
The conclusion of Theorem 2.4 follows from the boundedness of (2.15) and (2.16). Indeed, if $u\in H^{s,\varphi}_{p}(\Omega)$, then $Tu$ lies in the target space of the operator (2.15), i.e. $u$ satisfy the properties indicated in the second sentence of this theorem. Conversely, if a distribution $u\in\mathcal{S}^{\prime}(\Omega)$ satisfy these properties, then $u=KTu\in H^{s,\varphi}_{p}(\Omega)$. Formula (2.14) is due to
@@
Considering $H^{s,\varphi}_{p}(\Omega)$ as a solution space of elliptic boundary-value problems, we need the space of the traces of all distributions $u\in H^{s,\varphi}_{p}(\Omega)$ on $\Gamma$, with $s>1/p$. The trace space is found to be the Besov space of generalized smoothness over $\Gamma$ with smoothness indexes $s-1/p$ and $\varphi$, as in the $\varphi(\cdot)\equiv 1$ case. It can be defined on the base of the relevant space over $\mathbb{R}^{n-1}$ with the help of local charts on $\Gamma$. Let us introduce the latter space.
Put $d:=n-1$. Choose a function $\beta_{0}\in C^{\infty}(\mathbb{R}^{d})$ such that $\beta_{0}(y)=1$ whenever $|y|\leq 1$ and that $\beta_{0}(y)=0$ whenever $|y|\geq 2$, with $y\in\mathbb{R}^{d}$. Assuming $1\leq k\in\mathbb{Z}$, we define a function $\beta_{k}(y):=\beta_{0}(2^{-k}y)-\beta_{0}(2^{1-k}y)$ of $y\in\mathbb{R}^{d}$. Thus, the system $\{\beta_{k}:0\leq k\in\mathbb{Z}\}$ is a $C^{\infty}$-resolution of unity on $\mathbb{R}^{d}$.
@@
The space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ belongs to the class of Besov spaces of generalized smoothness investigated, e.g., in [1, 9, 10, 14, 20, 35, 39]. This space does not depend up to equivalence of norms on our choice of the function $\beta_{0}$ [9, Subsection 2.4]. If $\varphi(\cdot)\equiv 1$, then $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is the Besov space $B^{s}_{p}(\mathbb{R}^{d})$ (denoted also by $B^{s}_{p,p}(\mathbb{R}^{d})$) with smoothness index $s$ and integral-exponent $p$ (and sum-exponent $p$). Every space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is complete and separable, is continuously embedded in $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, with $\mathcal{S}(\mathbb{R}^{d})$ being continuously and densely embedded in $B^{s,\varphi}_{p}(\mathbb{R}^{d})$. This follows from the analogous properties of the Besov space $B^{0}_{p}(\mathbb{R}^{d})$ because the mapping $w\mapsto w_{s,\varphi}$ (defined by (2.1)) sets a topological isomorphism between $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ and $B^{0}_{p}(\mathbb{R}^{d})$ (see, e.g., [39, Proposition 7]). Hence, $C^{\infty}_{0}(\mathbb{R}^{d})$ is dense in $B^{s,\varphi}_{p}(\mathbb{R}^{d})$.
@@
They follow directly from (2.17) and the fact that $c^{-1}t^{-\varepsilon}\leq\varphi(t)\leq c\,t^{\varepsilon}$ whenever $t\geq 1$ with some number $c\geq 1$, due to property (i) of the class $\Upsilon$ (see [58, Section 1.5, property $1^{\circ}$]). The spaces $B^{s,\varphi}_{p}(\mathbb{R}^{d})$, where $s\in\mathbb{R}$ and $\varphi\in\Upsilon$, form the refined Besov scale (over $\mathbb{R}^{d}$) with integral-exponent $p$. Note that the Hilbert spaces $B^{s,\varphi}_{2}(\mathbb{R}^{d})$ and $H^{s,\varphi}_{2}(\mathbb{R}^{d})$ coincide up to equivalence of norms [39, Subsection 4.3, formula (19)].
@@
If $\varphi(\cdot)\equiv 1$, then $B^{s,\varphi}_{p}(\Gamma)$ is the Besov space $B^{s}_{p}(\Gamma)$ (see, e.g, [62, Definition 3.6.1]). Generally, we have the continuous embeddings (2.18) with $\Gamma$ instead of $\mathbb{R}^{d}$. They are compact due to the fact that $\ell<s$ implies the compact embedding $B^{s}_{p}(\Gamma)\hookrightarrow B^{\ell}_{p}(\Gamma)$ of Besov spaces. Hence, $B^{s,\varphi}_{p}(\Gamma)$ is continuously embedded in $\mathcal{D}^{\prime}(\Gamma)$.
@@
$[B^{s_{0},\varphi_{0}}_{p_{0}}(W),B^{s_{1},\varphi_{1}}_{p_{1}}(W)]_{\theta}=B^{s,\varphi}_{p}(W)$ (2.21)

### v2

@@ Abstract.
We build a solvability theory of elliptic boundary-value problems in normed Sobolev spaces of generalized smoothness for any integrability exponent $p>1$. The smoothness is given by a number parameter and a supplementary function parameter that varies slowly at infinity. These spaces are obtained by a combination of the methods of the complex interpolation with number parameter between Banach spaces and the quadratic interpolation with function parameter between Hilbert spaces applied to classical Sobolev spaces. We show that the spaces under study admit localization near a smooth boundary and describe their trace spaces in terms of Besov spaces with the same supplementary function parameter. We prove that a general differential elliptic problem induces Fredholm bounded operators on appropriate pairs of the spaces under study. We also find exact sufficient conditions for solutions of the problem to have a prescribed generalized or classical smoothness on a given set and establish corresponding a priori estimates of the solution. These results are specified for parameter-elliptic problems.
@@
Sobolev spaces; Besov spaces; generalized smoothness; interpolation functor; elliptic problem; parameter-elliptic problem; Fredholm operator; a priori estimate
@@ 1. Introduction
In the last two decades, the mathematical literature on function spaces demonstrates a growing interest in distribution spaces of generalized smoothness and their applications (see, e.g., [1, 7, 10, 11, 14, 21, 22, 35, 48, 49, 50]). These spaces characterize the smoothness properties of distributions by means of a function parameter depending on the frequency variables (or by a number sequence representing this function). This allows describing such properties far more finely than the scales of classical distribution spaces, which proved to be very useful in approximation theory [59], theory of stochastic process [25], and theory of partial differential equations (PDEs) [44, 52]. As to boundary-value problems for PDEs, the applications of distribution spaces of generalized smoothness has been restricted only to Hilbert spaces [5, 6, 36, 37, 43, 44], which can be investigated with the help of Spectral Theorem applied to a self-adjoint operator (see, e.g., [48]). However, theorems on solvability of various classes of boundary-value problems are well known for non-Hilbert distribution spaces of the Hölder or Sobolev smoothness given by a number (see, e.g., [2, 34, 38, 61, 62]). Concerning elliptic boundary-value problems, the expediency and prospects of their study in normed spaces of generalized smoothness were discussed by Triebel in [63, pp. 57–60].
@@
The purpose of this paper is to extend the classical theory of solvability of elliptic boundary-value problems in $L_{p}$-Sobolev spaces $H^{s}_{p}$ with $1<p<\infty$ (see, e.g., [2, Chapter V] and [62, Chapter 5]) to some Sobolev spaces of generalized smoothness. The latter spaces form the scale $\{H^{s,\varphi}_{p}\}$ calibrated with the help of a supplementary function parameter $\varphi$ much more finely than the Sobolev scale. The positive function $\varphi(t)$ of $t\geq 1$ varies slowly at infinity in the sense of Karamata, which implies that the space $H^{s,\varphi}_{p}$ is located between any spaces $H^{s-\varepsilon}_{p}$ and $H^{s+\varepsilon}_{p}$ with $\varepsilon>0$ on this scale. We study interpolation properties of the scale and show that every space $H^{s,\varphi}_{p}$ is obtained by the complex interpolation between certain spaces $H^{s_{0},\varphi_{0}}_{2}$ and $H^{s_{1}}_{p_{1}}$. Since the scale $\{H^{s,\varphi}_{2}\}$ of Hilbert spaces is obtained by the quadratic interpolation with function parameter between inner product Sobolev spaces [44, Theorem 1.14], we use these two interpolation methods to study the spaces $H^{s,\varphi}_{p}$ and elliptic problems in them. It is important for applications and follows from these interpolation properties that the space $H^{s,\varphi}_{p}(\Omega)$ over a bounded Euclidean domain $\Omega$ with a boundary $\Gamma\in C^{\infty}$ admits localization and that the trace operator acts continuously from whole $H^{s,\varphi}_{p}(\Omega)$ onto the Besov space $B^{s-1/p,\varphi}_{p}(\Gamma)$ of the same supplementary generalized smoothness $\varphi$ whenever $s>1/p$. The distribution space $B^{\sigma,\varphi}_{p}(\Gamma)$ with arbitrary $\sigma\in\mathbb{R}$ also admits localization and has quite similar interpolation properties. These results are discussed in Section 2 of the paper.
@@
Section 4 gives an application of spaces (1.1) to a general parameter-elliptic boundary-value problem. We show that this problem sets an isomorphism between these spaces for sufficiently large absolute values of the parameter from the ellipticity angle of the problem. Moreover, the solutions of the problem admit a corresponding two-sided a priori estimate in parameter-dependent norms with constants that do not depend on the parameter. This result is based on interpolation properties of such norms. The results can be applied to the study of parabolic initial-boundary-value problems in some anisotropic distribution spaces of generalized smoothness (cf. [4, Chapter II]).
@@ 2. Spaces of generalized smoothness
belongs to $L_{p}(\mathbb{R}^{n})$. Here, as usual, $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is the linear topological space of all tempered distributions on $\mathbb{R}^{n}$; thus, $\mathcal{S}^{\prime}(\mathbb{R}^{n})$ is dual to the linear topological Schwartz space $\mathcal{S}(\mathbb{R}^{n})$ of all infinitely smooth rapidly decreasing functions on $\mathbb{R}^{n}$. Moreover, $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the Fourier transform and its inverse; $\langle\xi\rangle:=(1+|\xi|^{2})^{1/2}$ is the smoothed absolute value of the vector $\xi\in\mathbb{R}^{n}$, and $L_{p}(\mathbb{R}^{n})$ is the normed space of all Lebesgue measurable functions $h:\mathbb{R}^{n}\to\mathbb{C}$ such that $|h|^{p}$ is integrable over $\mathbb{R}^{n}$. The space $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ is endowed with the norm
@@
The space $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ belongs to the classes of Sobolev spaces of generalized smoothness introduced and investigated by Volevich and Paneach [64, § 13], Schechter [57, Section 4], Goldman [17], Triebel [60, Section 5], Merucci [39, Section 4], and others. This space is complete and separable and is continuously embedded in $\mathcal{S}^{\prime}(\mathbb{R}^{n})$, with $\mathcal{S}(\mathbb{R}^{n})$ being continuously and densely embedded in $H^{s,\varphi}_{p}(\mathbb{R}^{n})$. This follows from the analogous properties of $L_{p}(\mathbb{R}^{n})$ in view of the evident fact that the mapping $w\mapsto w_{s,\varphi}$ sets an isometric isomorphism between $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ and $L_{p}(\mathbb{R}^{n})$. Hence, the set $C^{\infty}_{0}(\mathbb{R}^{n})$ of infinitely smooth compactly supported functions on $\mathbb{R}^{n}$ is dense in $H^{s,\varphi}_{p}(\mathbb{R}^{n})$.
If $\varphi(t)=1$ whenever $t\geq 1$, then $H^{s,\varphi}_{p}(\mathbb{R}^{n})$ is the Sobolev space $H^{s}_{p}(\mathbb{R}^{n})$ with smoothness index $s$ and integrability exponent $p$. Generally, we have the dense continuous embeddings
@@
Indeed, it follows from property (i) that $t^{-\varepsilon/2}\varphi(t)\to 0$ and $t^{\varepsilon/2}\varphi(t)\to\infty$ as $t\to\infty$ [58, Section 1.5, property $1^{\circ}$]; hence, the functions $t^{s}\varphi(t)/t^{s+\varepsilon}$ and $t^{s-\varepsilon}/t^{s}\varphi(t)$ of $t\geq 1$ are integrable over $1,\infty)$ with respect to $dt/t$, which implies (2.2) by [39, Subsection 4.1, Proposition 6].
@@
is naturally called the refined Sobolev scale (over $\mathbb{R}^{n}$) with integrability exponent $p$. In the $p=2$ case of Hilbert spaces, such a scale was introduced by Mikhailets and Murach [40] and applied to elliptic operators and elliptic boundary-value problems [41, 42, 43, 44, 51]. Various distribution spaces of generalized smoothness given by a slowly varying function were investigated by Haroske and Moura [20], specifically Besov spaces over $\mathbb{R}^{n-1}$ used below.
@@
$\{w\in H^{s,\varphi}_{p}(\mathbb{R}^{n}):\mathrm{supp}\,w\subset\mathbb{R}^{n}\setminus\Omega\}.$
@@
If $p\neq 2$, then the refined Sobolev scale with integrability exponent $p$ is obtained by complex interpolation between Hilbert spaces $H^{s_{0},\varphi_{0}}_{2}(\Omega)$ and Sobolev spaces $H^{s_{1}}_{p_{1}}(\Omega)$. Moreover, the collection of these scales is stable with respect to this interpolation. Let us formulate these properties in the form of two theorems. As usual, $E_{0},E_{1}]_{\theta}$ denotes the result of the complex interpolation with the number parameter $\theta\in(0,1)$ between Banach spaces $E_{0}$ and $E_{1}$ continuously embedded in a Hausdorff topological linear space (see., e.g., [62, Section 1.9]). Such a pair of the spaces is called an interpolation pair.
@@
$H^{s_{0},\varphi_{0}}_{p_{0}}(\Omega),H^{s_{1},\varphi_{1}}_{p_{1}}(\Omega)]_{\theta}=H^{s,\varphi}_{p}(\Omega)$ (2.9)
@@
Let $s\in\mathbb{R}$, $\varphi\in\Upsilon$, and $p\in(1,\infty)$ with $p\neq 2$. We arbitrarily chose a real number $p_{1}$ such that $1<p_{1}<p<2$ or $2<p<p_{1}<\infty$ and define a number $\theta\in(0,1)$ by the formula
@@
Note that the interpolation formulas (2.7), (2.9), and (2.11) are especially useful when we study bounded linear operators on the corresponding spaces because the boundedness of such operators is preserved under the interpolation between these spaces. Namely, if $E_{0},E_{1}]$ and $G_{0},G_{1}]$ are interpolation pairs of Banach spaces and if a linear mapping $T$ is a bounded operator from $E_{j}$ to $G_{j}$ for each $j\in\{0,1\}$, then $T$ is a bounded operator from $E_{0},E_{1}]_{\theta}$ to $G_{0},G_{1}]_{\theta}$ whenever $0<\theta<1$. The quadratic interpolation between Hilbert spaces has an analogous property.
@@
$\mathbb{R}^{n}_{+}:=\{(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}>0\},$
@@
$\partial\mathbb{R}^{n}_{+}=\mathbb{R}^{n}_{0}:=\{(x^{\prime},x_{n}):x^{\prime}\in\mathbb{R}^{n-1},x_{n}=0\in\mathbb{R}\}.$
@@
Choose functions $\chi_{1},\ldots,\chi_{\varkappa+r}\in C^{\infty}(\mathbb{R}^{n})$ such that $\mathrm{supp}\,\chi_{j}\subset U_{j}$ whenever $1\leq j\leq\varkappa+r$ and that $\chi_{1}+\cdots+\chi_{\varkappa+r}=1$ on $\overline{\Omega}$. Let $\mathcal{O}$ and $\mathcal{O_{+}}$ denote the operators of the extension of a function/distribution by zero onto $\mathbb{R}^{n}$ and $\mathbb{R}^{n}_{+}$, resp.
@@
where $u\in\mathcal{S}\,^{\prime}(\Omega)$, induces a bounded linear operator
@@
where $v_{1},\ldots,v_{\varkappa}\in\mathcal{S}^{\prime}(\mathbb{R}^{n}_{+})$ and $v_{\varkappa+1},\ldots,v_{\varkappa+r}\in\mathcal{S}^{\prime}(\mathbb{R}^{n})$. Here, $\mathcal{O}_{\Omega}$ denotes the operator of the extension of a distribution by zero onto $\Omega$. Then $KTu=u$ for every $u\in\mathcal{S}^{\prime}(\Omega)$. This mapping induces a bounded linear operator
@@
The conclusion of Theorem 2.4 follows from the boundedness of (2.15) and (2.16). Indeed, if $u\in H^{s,\varphi}_{p}(\Omega)$, then $Tu$ lies in the target space of the operator (2.15), i.e. $u$ satisfies the properties indicated in the second sentence of this theorem. Conversely, if a distribution $u\in\mathcal{S}^{\prime}(\Omega)$ satisfies these properties, then $u=KTu\in H^{s,\varphi}_{p}(\Omega)$. Formula (2.14) is due to
@@
Considering $H^{s,\varphi}_{p}(\Omega)$ as a solution space of elliptic boundary-value problems, we need the space of the traces of all distributions $u\in H^{s,\varphi}_{p}(\Omega)$ on $\Gamma$, with $s>1/p$. The trace space is found to be the Besov space of generalized smoothness over $\Gamma$ with smoothness indexes s1/ps-1/p and $\varphi$, as in the $\varphi(\cdot)\equiv 1$ case. It can be defined on the base of the relevant space over $\mathbb{R}^{n-1}$ with the help of local charts on $\Gamma$. Let us introduce the latter space.
Put d:=n1d:=n-1. Choose a function $\beta_{0}\in C^{\infty}(\mathbb{R}^{d})$ such that $\beta_{0}(y)=1$ whenever $|y|\leq 1$ and that $\beta_{0}(y)=0$ whenever $|y|\geq 2$, with $y\in\mathbb{R}^{d}$. Assuming $1\leq k\in\mathbb{Z}$, we define a function $\beta_{k}(y):=\beta_{0}(2^{-k}y)-\beta_{0}(2^{1-k}y)$ of $y\in\mathbb{R}^{d}$. Thus, the system $\{\beta_{k}:0\leq k\in\mathbb{Z}\}$ is a $C^{\infty}$-resolution of unity on $\mathbb{R}^{d}$.
@@
The space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ belongs to the class of Besov spaces of generalized smoothness investigated, e.g., in [1, 9, 10, 14, 20, 35, 39]. This space does not depend up to equivalence of norms on our choice of the function $\beta_{0}$ [9, Subsection 2.4]. If $\varphi(\cdot)\equiv 1$, then $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is the Besov space $B^{s}_{p}(\mathbb{R}^{d})$ (denoted also by $B^{s}_{p,p}(\mathbb{R}^{d})$) with smoothness index $s$ and integrability exponent $p$ (and summability exponent $p$). Every space $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ is complete and separable, is continuously embedded in $\mathcal{S}^{\prime}(\mathbb{R}^{d})$, with $\mathcal{S}(\mathbb{R}^{d})$ being continuously and densely embedded in $B^{s,\varphi}_{p}(\mathbb{R}^{d})$. This follows from the analogous properties of the Besov space $B^{0}_{p}(\mathbb{R}^{d})$ because the mapping $w\mapsto w_{s,\varphi}$ (defined by (2.1)) sets a topological isomorphism between $B^{s,\varphi}_{p}(\mathbb{R}^{d})$ and $B^{0}_{p}(\mathbb{R}^{d})$ (see, e.g., [39, Proposition 7]). Hence, $C^{\infty}_{0}(\mathbb{R}^{d})$ is dense in $B^{s,\varphi}_{p}(\mathbb{R}^{d})$.
@@
They follow directly from (2.17) and the fact that $c^{-1}t^{-\varepsilon}\leq\varphi(t)\leq c\,t^{\varepsilon}$ whenever $t\geq 1$ with some number $c\geq 1$, due to property (i) of the class $\Upsilon$ (see [58, Section 1.5, property $1^{\circ}$]). The spaces $B^{s,\varphi}_{p}(\mathbb{R}^{d})$, where $s\in\mathbb{R}$ and $\varphi\in\Upsilon$, form the refined Besov scale (over $\mathbb{R}^{d}$) with integrability exponent $p$. Note that the Hilbert spaces $B^{s,\varphi}_{2}(\mathbb{R}^{d})$ and $H^{s,\varphi}_{2}(\mathbb{R}^{d})$ coincide up to equivalence of norms [39, Subsection 4.3, formula (19)].
@@
If $\varphi(\cdot)\equiv 1$, then $B^{s,\varphi}_{p}(\Gamma)$ is the Besov space $B^{s}_{p}(\Gamma)$ (see, e.g, [62, Definition 3.6.1]). Generally, we have the continuous embeddings (2.18) with $\Gamma$ instead of $\mathbb{R}^{d}$. They are compact due to the fact that $\ell<s$ implies the compact embedding $B^{s}_{p}(\Gamma)\hookrightarrow B^{\ell}_{p}(\Gamma)$ of Besov spaces. Moreover, $B^{s,\varphi}_{p}(\Gamma)$ is continuously embedded in $\mathcal{D}^{\prime}(\Gamma)$.
@@
$B^{s_{0},\varphi_{0}}_{p_{0}}(W),B^{s_{1},\varphi_{1}}_{p_{1}}(W)]_{\theta}=B^{s,\varphi}_{p}(W)$ (2.21)